OpenStudy (anonymous):

Can someone help me with some inequality problems, please? I'll give a medal! :D

OpenStudy (anonymous):

Mark and Nora both work at the same pet store. The inequality mc017-1.jpg relates x, the number of hours Mark works each week, and y, the number of hours that Nora works. Which ordered pair must be included in the solution set of the inequality when interpreted in the context of the problem? (–7, 20) (0, 0) (8, 2) (13, –28) (–7, 20) is one of the solutions, but (0, 0) is on the line. Which is correct?

OpenStudy (princeharryyy):

whers the fig

OpenStudy (princeharryyy):

@Catseyeglint911

OpenStudy (anonymous):

This is the graph of the equation.

OpenStudy (princeharryyy):

-7,20 isn't on the line @Catseyeglint911

OpenStudy (anonymous):

I know, but it is in the solution set (the shaded area). That's what's confusing to me.

OpenStudy (princeharryyy):

do one thing write the equation of line thats shown

OpenStudy (anonymous):

I'm so sorry. I just realized that the equation wasn't copied with the word problem. \[y \ge 2x\]

OpenStudy (princeharryyy):

r u sure @Catseyeglint911

OpenStudy (anonymous):

Am I sure that's the right equation? I am.

OpenStudy (princeharryyy):

thats gud !!! now what do u think about the answer

OpenStudy (princeharryyy):

now what are ur x and y's

OpenStudy (princeharryyy):

@Catseyeglint911

OpenStudy (princeharryyy):

x and y must be the hours and hours can't be negative ... i think

OpenStudy (anonymous):

I guess that (0,0) is correct since when you plug it into the equation, it's true, but so is (–7, 20)

OpenStudy (anonymous):

\[20 \ge 2(-7)\] \[20 \ge 14\] True

OpenStudy (princeharryyy):

so x can't be negative so as y

OpenStudy (princeharryyy):

can the hours be negative @Catseyeglint911

OpenStudy (anonymous):

So just to clarify...what you're saying is that even though the coordinates (–7, 20) holds true when plugged into the equation, it is still wrong because the hours can't be negative, correct?

OpenStudy (princeharryyy):

yeah!!!! the only solutions which are in the 1st coordinates are the solutions @Catseyeglint911

OpenStudy (anonymous):

Oh wow I feel retarded. Thank so much for your help. :) Could I bother you for one more question by chance?

OpenStudy (princeharryyy):

and 8,2 doesn't satisfy the equation so the solution is 0,0

OpenStudy (princeharryyy):

ohk !!!

OpenStudy (anonymous):

Lionel plays trumpet for a minimum of 45 minutes on the days that he practices. If x is the number of days that Lionel practices and y is the total number of hours he spends practicing, which inequality represents this situation? \[y \ge 0.75x\] \[0.75y \ge x\] \[y \le 0.75x\] \[0.75y \le x\]

OpenStudy (anonymous):

I pretty sure it's either the first one or the second one since 0.75 is the minimum amount of time he plays per day, and the x is the number of days.

OpenStudy (princeharryyy):

don't u think that 1st one's the ans

OpenStudy (princeharryyy):

wait a sec somethings confusing

OpenStudy (anonymous):

i hope this helps http://www.mathportal.org/calculators.php

OpenStudy (anonymous):

Sorry i meant the first or the third. I don't think the second or forth is right.

OpenStudy (princeharryyy):

see x>= y so now i'm sure u can catch up with the answer @Catseyeglint911

OpenStudy (anonymous):

Ah so it is the third

OpenStudy (princeharryyy):

yup !!!

OpenStudy (anonymous):

It said it was incorrect. D: I'm so confused.

OpenStudy (princeharryyy):

u r confusing me too .... i was talking to someone so wait....

OpenStudy (anonymous):

okay.

OpenStudy (princeharryyy):

day 1 => y=(45/60)*1 = 0.75*1; day 2 => y= 0.75*2 day 3 => y=0.75*3 now as 0.75 is the minimum hrs he can play imply he can play more then 0.75 as well imply the increase in y so, as 0.75 increases so y increases => y >= 0.75x option c @Catseyeglint911

OpenStudy (princeharryyy):

sorry this is option a @Catseyeglint911

OpenStudy (princeharryyy):

not c

OpenStudy (anonymous):

Oh okay, that makes sense.

OpenStudy (anonymous):

Thank you for your help again! :)

OpenStudy (princeharryyy):

u r welcome @Catseyeglint911

OpenStudy (anonymous):

what is the question

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

og I thought there were like 3 questions I didn't know all were solved lol

OpenStudy (anonymous):

Nope, but thank you for offering. :)

OpenStudy (princeharryyy):

@anonymous_secret thnks for ur medal and u too @Catseyeglint911

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

According to the online test, it is (0,0)